Circular No. 7492

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2000 R2 (LINEAR) An object reported as asteroidal by the LINEAR team on Sept. 3, and

subsequently posted on the NEO Confirmation Page, has been reported as cometary in appearance on various CCD exposures: Sept. 7.2 UT, condensed 9′′ coma (D. Balam, Victoria, 1.82-m reflector); Sept. 7.9, object seemed diffuse with a faint east-west tail ∼ 10′′ long (M. Tichý, Kleť, 0.57m reflector); Sept. 18.51, near-stellar condensation with a fairly narrow 40′′ tail in p.a. 90^{o} (R. H. McNaught, Siding Spring, 1.0-m reflector + 100s R frames). The available astrometric observations are given on MPEC 2000-S04, where orbital elements show this to be a short-period comet.

2000 UT α_{2000} δ_{2000} m_{2} Observer
Sept.3.14818 19^{h}54^{m}00.^{s}63 −14^{o}41′32′′.0 18.7 LINEAR
T = 2000 Sept.12.7665 TT  ω = 147.^{o}2011
e = 0.590057  Ω = 187.3825 }2000.0
q = 1.392287 AU  i = 3.2336
a = 3.396297 AU n^{o} = 0.1574692 P = 6.259 years

V723 CASSIOPEIAE D. K. Lynch, R. J. Rudy, S. Mazuk, and C. Venturini, The Aerospace

Corporation; R. C. Puetter, Center for Astrophysics and Space Science, University of California at San Diego; and T. Armstrong, Thomas K. Armstrong Development Corporation, report 0.8–2.5-μm spectrophotometry of V723 Cas = N Cas 1995 using the Lick Observatory 3-m telescope and the Aerospace Near Infrared Imaging Spectrograph (NIRIS) on July 20.45 UT, ∼ 4.6 yr after peak brightness: “The spectrum resembled that of a late-stage classical nova, displaying forbidden lines of singly and doubly ionized sulfur and neutral nitrogen, along with strong permitted lines of H I, He I, and He II. The strongest feature is He I 1.0830-μm. Very strong coronal lines are also present, including [Si VI] at 1.9629 μm, [Ca VIII] at 2.3214 μm, [Si VII] at 2.4827 μm, and possibly [Al IX] at 2.0469 μm. The underlying continuum decreased toward the red and was indicative of free-free emission. There was no evidence of thermal emission from dust in the region 1–2.5 μm. J, H, and K magnitudes determined from the spectrum are 12.6, 12.6, and 11.7, respectively. The shape of the spectrum seems to have changed little since the 1999 report by Rudy et al. (IAUC 7259), although its absolute level has decreased by about a factor of 2.”

2000 September 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green